c) A dolphin leaps out of the water at an angle of 36° above the horizontal. The horizontal component of the dolphin's velocity is 9.0 m/s. Find the magnitude of the vertical component of the velocity.

d.) A golf ball rolls off a horizontal cliff with an initial speed of 12.9 m/s. The ball falls a vertical distance of 15.5 m into a lake below.

How much time does the ball spend in the air?

Don't do this part, I know it is 1.777

What is the speed v of the ball just before it strikes the water?

This is the part I need.
Answer #1

C) V(Cos36⁰) = 9

V = 11.1246 m/s

Vertical velocity = 11.1246×Sin36⁰ = 6.54 m/s

D) t = sqrt(2h/g) = sqrt(2×15.5/9.8) = 1.778 s

Vy = 9.8 × 1.778 = 17.43 m/s

Vx = 12.9 m/s

V = sqrt (Vx² + Vy²) = sqrt ( 17.43² + 12.9²) = 21.68 m/s